OpenStudy (anonymous):

Compare 8 ⋅ 107 to 2 ⋅ 104. 8 ⋅ 107 is 4,000 times larger than 2 ⋅ 104. 8 ⋅ 107 is 400 times larger than 2 ⋅ 104. 8 ⋅ 107 is 40 times larger than 2 ⋅ 104. 8 ⋅ 107 is 4 times larger than 2 ⋅ 104.

OpenStudy (texaschic101):

8 * 107 = 856 2 * 104 = 208 856/208 = 4.11 solution : 8 * 107 is 4 times larger then 2 * 104